# Competing Effects of Local Solvation Structures on Chemical Shift Changes of Liquid Electrolyte

### v1

@@ 3.2 The Local Structures of the LiFSI/DME Electrolytes
The changes observed in the NMR spectra are a direct consequence of the evolving solvation structures. At the heart of this transformation lies the intricate interplay between cation-solvent and cation-anion interactions, which gives rise to distinct categories of the combinations of the cations, solvents and anions, such as SSIPs, CIPs, and AGGs, see Fig. 3(a). Before interpreting the complicated chemical shift variations due to the changes in LiFSI concentration, we first focus on the classifications of the solvation structures, based on the number of anions coordinated with Li+, as shown in Fig. 3(b), which demonstrates that the high-number FSI- category (more than two) dominates as the concentration increases. Not surprisingly, in the low concentration case (1 M), the solvated Li+/FSI- ions are separated by the solution DME molecules in the LiFSI/DME electrolytes, therefore, the coordination number of FSI- anions to Li+ ion is 0 or 1, which are corresponding to SSIP and CIP scenario. When the LiFSI concentration increases (2 or 3 M), the ratios of the SSIPs () and CIPs () decrease, the Li+ ions are more bonded with FSI- anions leading to the increase of the AGGs ( or 3). As more LiFSI salt is dissolved in the DME solution (4 M), approaching the saturation limit (5.3 M)37, the situation changes dramatically. The ratio of SSIPs and CIPs no longer continue to decrease but instead shows a slight increase compared to the 3 M case. Meanwhile, the proportion of lower-order AGGs ( or 3, noted as AGGs) surprisingly decreases, while the proportion of higher-order AGGs (, noted as AGGs+) continues to increase.
@@
The terminology of SSIPs, CIPs, and AGGs can provide a framework for quantifying the solvation structures surrounding Li+ ions. Meanwhile, the degree of inhomogeneity in the local molecular environment of Li+ ions plays a crucial role in determining electron density around Li+ nucleus, which in turn influences dynamic shielding effects and the resulting NMR chemical shift. Here, we utilized an order parameter, which associates local structure index (LSI) 61, 62, 63 to the individual Li+ ions and its local neighbor oxygen pairs. In short, the LSI order parameter is the mean-squared-deviation among the radial distances corresponding to the set of the oxygen atoms that surround a given Li+ ion, the LSI value is assigned as the inhomogeneity in the distribution of radial distances. The details of the definition can be referred to Sect. VI in SI. The schematic diagram provided in the insert of Fig. 3(c), which illustrates the different local atomic environments that can be distinguished and quantified by the LSI order parameter. For instance, a Li+ ion with a high LSI value is typically found in a more ordered local environment, where neighboring oxygen atoms are densely concentrated around 2.0 Å and sparsely distributed around 3.0 Å. This results in a clearer separation between the first and second coordination shells and a relatively low local atomic number density (as depicted on the right of the inset). Consequently, the electron density around the central Li+ ion is less perturbed by the surrounding oxygen atoms, potentially leading to lower (absolute) NMR chemical shift values. Conversely, a Li+ ion with a low LSI value resides in a locally disordered environment. This is characterized by a relatively high packing of neighboring oxygen atoms in the interstitial region and an elevated local atomic number density (as illustrated on the left of the inset). In such a densely packed and disordered atomic environment, the delicate balance of the surrounding electron density is disrupted, resulting in a higher NMR chemical shift value.
@@
Now, Let us focus on the impact of LiFSI concentration changes on the LSI distribution, as shown in Fig. 3(c). At low concentrations, the LSI distribution exhibits a broad profile, suggesting a relatively ordered packing of neighboring oxygen atoms in the interstitial region. As the concentration of LiFSI increases in the DME solution, the interstitial region around the central Li+ ion becomes more compressed. Consequently, the LSI distribution sharpens and displays a more pronounced peak. Consequently, the LSI distribution sharpens and displays a more pronounced peak. Simultaneously, a shoulder emerges at LSI = 0.15, indicating a concentration of neighboring oxygen atoms around 2.0, which forms a distinct sub-interstitial structure. As there are much more oxygen atoms aggregating around Li+ ions, the sub-interstitial occurs between the first and the second coordination shell, which can be corroborated by the forward shift of the second coordination shell at 4 M, as shown in Fig. S3(c). This structure can also be analyzed using the RDFs between the Li+ ions, see Fig. S3(d) of the SI. Notably, in the 4 M case, an additional peak appears around 3 Å in the Li+-Li+ RDF, indicating a direct interaction between two Li+ ions. This suggests that the solvation structures of neighboring Li+ ions are merging. Consequently, locally high-density AGGs+-like structures, also referred to as micelle-like structures 13, 72, begin to emerge, as illustrated schematically in Fig. S6 of the SI. Under these conditions, the remaining Li+ ions have a higher probability of forming ion pairs, such as SSIPs and CIPs. As shown in Fig. 3(b), the ratio of SSIPs and CIPs no longer decreases but instead exhibits a slight increase compared to the 3 M case.
@@ 3.3 Connections between Local Structures and NMR Chemical Shift
Notably, through structural-chemical shift analysis, we categorize the data points based on their distinct NMR values. The representative species corresponding to different NMR values exhibit strong correlations with their local chemical environments. Specifically, points with the lowest NMR values (red) are typically associated with SSIPs (), while those with intermediate NMR values (yellow) correspond to CIPs (). The dominant chemical shift region (green) is linked to AGGs ( or 3), and the highest NMR values are observed for AGGs+ (). The detailed assignment can be found in the Sect. VII of the SI. This finding aligns with chemical intuition, as discussed in the previous section, and underscores how the unsupervised PCA method effectively captures the influence of variations in the local environment on surrounding electron density.
@@
Now, we can further investigate the changes in their populations to the varying LiFSI concentrations, as shown in Fig. 4(b). Notably, a dramatic shift occurs between 3 M and 4 M. Clearly, the most dominant structures are those with intermediate NMR values (yellow) and those with higher NMR values (green). The former shows a sudden increase, while the latter experiences a sharp decrease. Building on the insights from the discussion from the previous section, we infer that the increase in CIP structures (yellow) is due to the emergence of highly localized AGGs+ structures induced by Li+-Li+ interactions. Moreover, the detailed decomposition of the PCA analysis with respect to the different LiFSI concentrations, shown in Fig. S7 of the SI, reveals similar distributions across the various concentrations. It clearly demonstrates the emergence of the CIP structures as the concentration increases from 3 M to 4 M.
@@
Going further, the results from the PCA are supported by the local structural parameters, such as the LSI. In Fig. 5, we present the NMR chemical shift values as a function of various LSI values. As the chemical shift values become higher, the LSI values simultaneously reduce, the tendency of which is consistent in four concentrations. According to the corresponding structures we assigned above, the average LSI value of blue region () at 4 M smaller than that at 3M, which means that the high hierarchy coordination number structures () replace this region instead of . The reason for this phenomenon is that the high hierarchy coordination number configurations tend to have a filled sub-interstitial and form several long chain clusters of Li+ and FSI- connected, the LSI of which are lower than those of clusters with . Moreover, the clusters with (yellow region) emerge along with the case at 4 M, surrounding these AGGs+-like structures, ultimately leading to the lower LSI compared with 3 M case. This trend aligns with the findings from the local structure analysis in the previous section. Thus, the trend highlights the complex interplay between different local environments, with notable shifts driven by concentration effects. Moreover, the highly-localized AGGs+ structures induced by Li+-Li+ interactions can further reduce the diffusion constants of Li+ ions, as they form additional ion pairs.73 A similar scenario is observed in the diffusion behavior of hydronium and hydroxide in water 74, the difference of the diffusion constant due to the presence of a hypercoordinate solvation structure. Nevertheless, the downfield shift of the 7Li NMR chemical shift from 3 M to 4 M, is the signature of the emergence of the highly-localized AGGs structures and resulting less coordinated CIPs structures in the high concentration electrolyte.
@@ 5.1 Data Availability Statement
The code of automated workflow (ai2-kit) can be found at https://github.com/chenggroup/ai2-kit, and datasets for NN can be found at xxx upon acceptance of the manuscript.
@@ 5.3 Notes
We thank Prof. Hai-ming Liu at ShanghaiTech University for helpful discussions. F. T. acknowledges the National Key R&D Program of China (Grant No. 2024YFA1210804) and a startup fund at Xiamen University. J. C. acknowledges the National Natural Science Foundation of China (Grant Nos. 22225302, 21991151, 21991150, 22021001, 92161113, and 20720220009) and the Laboratory of AI for Electrochemistry (AI4EC) and IKKEM (Grant Nos. RD2023100101 and RD2022070501) for financial support. This work used the computational resources in the IKKEM intelligent computing center.

### v2

@@ 3.2 Local Structures of the LiFSI/DME Electrolytes
When increasing the concentration, the changes observed in the NMR spectra are a direct consequence of the evolving solvation structures. At the heart of this transformation lies the intricate interplay between cation-solvent and cation-anion interactions, which gives rise to distinct categories of the combinations of the cations, solvents and anions, such as SSIPs, CIPs, and AGGs, see Fig. 3(a). Before interpreting the complicated chemical shift variations due to the changes in LiFSI concentration, we first focus on the classifications of the solvation structures, based on the number of anions coordinated with Li+, as shown in Fig. 3(b), which demonstrates that the high-number FSI- category (more than two) dominates as the concentration increases. Not surprisingly, in the low concentration case (1 M), the solvated Li+/FSI- ions are separated by the solution DME molecules in the LiFSI/DME electrolytes, therefore, the coordination number of FSI- anions to Li+ ion is 0 or 1, which are corresponding to SSIP and CIP scenario. When the LiFSI concentration increases (2 or 3 M), the ratios of the SSIPs () and CIPs () decrease, and the Li+ ions are more bonded with FSI- anions leading to the increase of the AGGs ( or 3). As more LiFSI salt is dissolved in the DME solution (4 M), approaching the saturation limit (5.3 M)37, the situation changes dramatically. The ratio of SSIPs and CIPs no longer continue to decrease but instead shows a slight increase compared to the 3 M case. Meanwhile, the proportion of lower-order AGGs ( or 3, noted as AGGs) surprisingly decreases, while the proportion of higher-order AGGs (, noted as AGGs+) continues to increase.
@@
The terminology of SSIPs, CIPs, and AGGs can provide a framework for quantifying the solvation structures surrounding Li+ ions. Meanwhile, the degree of inhomogeneity in the local molecular environment of Li+ ions plays a crucial role in determining electron density around the Li+ nucleus, which in turn influences dynamic shielding effects and the resulting NMR chemical shift. Here, we utilized an order parameter, which associates local structure index (LSI)61, 62, 63 to the individual Li+ ions and their local neighboring oxygen pairs. In short, the LSI order parameter is the mean-squared-deviation among the radial distances corresponding to the set of the oxygen atoms that surround a given Li+ ion, and the LSI value is assigned as the inhomogeneity in the distribution of radial distances. The detailed definition can be referred to Sect. VII in SI. The schematic diagram provided in the insert of Fig. 3(c) illustrates the different local atomic environments that can be distinguished and quantified by the LSI order parameter. For instance, a Li+ ion with a high LSI value is typically found in a more ordered local environment, where neighboring oxygen atoms are densely concentrated around 2.0 Å and sparsely distributed around 3.0 Å. This results in a clearer separation between the first and second coordination shells and a relatively low local atomic number density (as depicted on the right of the inset). Consequently, the electron density around the central Li+ ion is less perturbed by the surrounding oxygen atoms, potentially leading to lower (absolute) NMR chemical shift values. Conversely, a Li+ ion with a low LSI value resides in a locally disordered environment. This is characterized by a relatively high packing of neighboring oxygen atoms in the interstitial region and an elevated local atomic number density (as illustrated on the left of the inset). In such a densely packed and disordered atomic environment, the delicate balance of the surrounding electron density is disrupted, resulting in a higher NMR chemical shift value.
@@
Now, let us focus on the impact of LiFSI concentration changes on the LSI distribution, as shown in Fig. 3(c). At low concentrations, the LSI distribution exhibits a broad profile, suggesting a relatively ordered packing of neighboring oxygen atoms in the interstitial region. As the concentration of LiFSI increases in the DME solution, the interstitial region around the central Li+ ion becomes more compressed. Consequently, the LSI distribution sharpens and displays a more pronounced peak. The distance distribution of the peak in 4 M spans from 1.8 Å to 3.0 Å with a bin size of 0.1 Å. However, the distance distribution of the shoulder, which emerges at LSI = 0.15 , is concentrated between 1.8 Å and 2.2 Å, then vanished before reappearing beyond 3.0 Å. The comparison means that the smaller the LSI value is, the tighter the first solvation shell is. As many more oxygen atoms are aggregating around Li+ ions, distinct sub-interstitial structures occur between the first and the second coordination shell, which can be corroborated by the forward shift of the second coordination shell at 4 M, as shown in Fig. S3(c). This structure can also be analyzed using the RDFs between the Li+ ions, see Fig. S3(d) of the SI. Notably, in the 4 M case, an additional peak appears around 3.0 Å in the Li+-Li+ RDF, indicating a direct interaction between two Li+ ions. This suggests that the solvation structures of neighboring Li+ ions are merging. Consequently, locally high-density AGGs+-like structures, also referred to as micelle-like structures13, 74, begin to emerge, as illustrated in Fig. 5(b) and (c). Under these conditions, the remaining Li+ ions have a higher probability of forming ion pairs, such as SSIPs and CIPs. As shown in Fig. 3(b), the ratio of SSIPs and CIPs no longer decreases but instead exhibits a slight increase compared to the 3 M case.
@@ 3.3 Connections between Local Structures and NMR Chemical Shift
Specifically, points with the lowest NMR values (red) are typically associated with SSIPs (), while those with intermediate NMR values (yellow) correspond to CIPs (). The dominant chemical shift region (green) is linked to AGGs ( or 3), and the highest NMR values are observed for AGGs () and AGGs+ (). The detailed assignment can be found in Tab. S4 of the SI. This finding aligns with chemical intuition, as discussed in the previous section, and underscores how the unsupervised PCA method effectively captures the influence of variations in the local environment on surrounding electron density.
@@
Now, we can further investigate the changes in their populations to the varying LiFSI concentrations, as shown in Fig. 4(b). Notably, a dramatic shift occurs between 3 M and 4 M. Clearly, the most dominant structures are those with intermediate NMR values (yellow) and those with higher NMR values (green). The former shows a sudden increase, while the latter experiences a sharp decrease. Building on the insights from the discussion from the previous section, we infer that the rise in CIP structures (yellow) is due to the emergence of highly localized AGGs+ structures induced by Li+-Li+ interactions. Moreover, the detailed decomposition of the PCA analysis for the different LiFSI concentrations, as shown in Fig. S7 of the SI, reveals similar distributions across the various concentrations. It clearly demonstrates the increased population of the CIPs as the concentration increases from 3 M to 4 M.
@@
Going further, the results from the PCA are supported by the local structural parameters, such as the n and LSI. In Fig. 5(a), we present the NMR chemical shift values as a function of various LSI values, with the color assignment according to n. The given Li+ ion in a higher n solvation environment tends to move towards the upfield direction in chemical shift, and vice versa, which provides evidence for the region assignment of solvation structures in the PCA. As the chemical shift values become higher, the LSI values simultaneously reduce, the tendency of which is consistent in four concentrations. The reason for this phenomenon is that the higher-order coordination number () configurations tend to have a filled sub-interstitial and form several long-chain clusters of Li+ and FSI- connected, the LSI of which are lower than those of clusters with . Moreover, the clusters with or 1 (blue region) emerge along with the case at 4 M, as shown in Fig. 5(d) and (e), surrounding these AGGs+-like structures, ultimately leading to the lower LSI compared with 3 M case.
The long-chain clusters shown in Fig. 5(d) and (e) tend to introduce the phase separation of DME and FSI-. We have analyzed the surrounding molecules around the central Li+ ions within 3.3 Å, using the midpoint of each molecule for measurement. The curves illustrated in Fig. 6 represent the probability density of the molecule appearance at a particular distance from the central Li+ ions. They are similar in 1 to 3 M for both DME and FSI-, where the main peak of FSI- appears at 3.15 Å, and DME shows a shoulder around 3.15 Å. In 4 M solutions, however, more FSI- anions penetrate the first solvation shell at 2.15 Å, and DME molecules aggregate 2.80 Å with the shoulder disappearing. The aggregation of FSI- ions in the inner shell and the clustering of DME molecules create distinct phases in solutions, resulting in localized molecular redistribution and aggregation. The corresponding schematic diagrams are shown in Fig. S8 of the SI. This trend aligns with the findings from the local structure analysis in the previous section, which highlights the complex interplay between different local environments, with notable shifts driven by concentration effects. Simultaneously, the ELF effects arise due to the strong electrostatic interactions and localized electron density around Li+ ions, which restricts the free movement of charge carriers. Meanwhile, the steric effects hinder the spatial pathways for ion migrations. These combined effects lead to a high viscosity and low conductivity in HCEs.78 Nonetheless, the distances between Li+ ions are much closer in 4 M than the other three concentrations, reducing the distance required for the transportation of Li+ ions to maintain the flux when Li+ ions near the electrode are depleted at high current density. Also, the solid electrolyte interphase (SEI) formed in 4 M LiFSI/DME shows a slow anion degradation during extended cycling, resulting in a substantial accumulation of inorganic components.79 These components enhance ionic conductivity and provide mechanical protection, strengthening the SEI and significantly improving the CE when using a Li metal anode.79, 80 Nevertheless, the downfield shift of the 7Li NMR chemical shift from 3 M to 4 M, is the signature of the emergence of the highly localized AGGs+ structures and results in less coordinated CIPs structures in the high-concentration electrolytes.
@@ 5.1 Data Availability Statement
[…]
@@ 5.2 Supporting Information
5.3 Author Contributions
#Q. Y. and Y. S. contributed equally to this work.
@@ 5.4 Notes
We thank Prof. Bing-wen Hu at East China Normal University and Prof. Hai-ming Liu at ShanghaiTech University for helpful discussions. F. T. acknowledges the National Key R&D Program of China (Grant No. 2024YFA1210804) and a startup fund at Xiamen University. J. C. acknowledges the National Natural Science Foundation of China (Grant Nos. 22021001, 22225302, 21991151, 21991150, 92161113, and 20720220009) and the Laboratory of AI for Electrochemistry (AI4EC) and IKKEM (Grant Nos. RD2023100101 and RD2022070501) for financial support. This work used the computational resources in the IKKEM intelligent computing center.